# Knowledge Graph Prompting for Multi-Document Question Answering

### 17 Dec 2023 v2

@@ Abstract
The ‘pre-train, prompt, predict’ paradigm of large language models (LLMs) has achieved remarkable success in open-domain question answering (OD-QA). However, few works explore this paradigm in the scenario of multi-document question answering (MD-QA), a task demanding a thorough understanding of the logical associations among the contents and structures of different documents. To fill this crucial gap, we propose a Knowledge Graph Prompting (KGP) method to formulate the right context in prompting LLMs for MD-QA, which consists of a graph construction module and a graph traversal module. For graph construction, we create a knowledge graph (KG) over multiple documents with nodes symbolizing passages or document structures (e.g., pages/tables), and edges denoting the semantic/lexical similarity between passages or intra-document structural relations. For graph traversal, we design an LLM-based graph traversal agent that navigates across nodes and gathers supporting passages assisting LLMs in MD-QA. The constructed graph serves as the global ruler that regulates the transitional space among passages and reduces retrieval latency. Concurrently, the graph traversal agent acts as a local navigator that gathers pertinent context to progressively approach the question and guarantee retrieval quality. Extensive experiments underscore the efficacy of KGP for MD-QA, signifying the potential of leveraging graphs in enhancing the prompt design for LLMs. Our code: https://github.com/YuWVandy/KG-LLM-MDQA.
@@ 4                   Terminate
  Method HotpotQA IIRC 2WikiMQA MuSiQue Comp Rank
Acc EM F1 Acc EM F1 Acc EM F1 Acc EM F1 Struct-EM w Comp w/o Comp
@@ 5.1 Performance Comparison on MD-QA
Moreover, without any particular design for document structures, no existing baselines can handle structural questions in Comp, e.g. ‘What is the difference between Page 1 and Page 2’ or ‘In Table 3, which station has the highest average flow rate?’. Fortunately, with the constructed KG incorporating structural nodes and our designed traversal algorithm retrieving structural contexts, our proposed method achieves 67% Struct-EM.
@@ Acknowledgements
This research is supported by Adobe Research and National Science Foundation (NSF) under grant number IIS2239881.

### 25 Dec 2023 v3

@@ Abstract
The ‘pre-train, prompt, predict’ paradigm of large language models (LLMs) has achieved remarkable success in open-domain question answering (OD-QA). However, few works explore this paradigm in multi-document question answering (MD-QA), a task demanding a thorough understanding of the logical associations among the contents and structures of documents. To fill this crucial gap, we propose a Knowledge Graph Prompting (KGP) method to formulate the right context in prompting LLMs for MD-QA, which consists of a graph construction module and a graph traversal module. For graph construction, we create a knowledge graph (KG) over multiple documents with nodes symbolizing passages or document structures (e.g., pages/tables), and edges denoting the semantic/lexical similarity between passages or document structural relations. For graph traversal, we design an LLM-based graph traversal agent that navigates across nodes and gathers supporting passages assisting LLMs in MD-QA. The constructed graph serves as the global ruler that regulates the transitional space among passages and reduces retrieval latency. Concurrently, the graph traversal agent acts as a local navigator that gathers pertinent context to progressively approach the question and guarantee retrieval quality. Extensive experiments underscore the efficacy of KGP for MD-QA, signifying the potential of leveraging graphs in enhancing the prompt design and retrieval augmented generation for LLMs. Our code: https://github.com/YuWVandy/KG-LLM-MDQA.
@@ 4                   Terminate
  Method HotpotQA IIRC 2WikiMQA MuSiQue PDFTriage Rank
Acc EM F1 Acc EM F1 Acc EM F1 Acc EM F1 Struct-EM w PDFTriage w/o PDFTriage
@@ 5.1 Performance Comparison on MD-QA
Moreover, without any particular design for document structures, no existing baselines can handle structural questions in PDFTriage, e.g. ‘What is the difference between Page 1 and Page 2’ or ‘In Table 3, which station has the highest average flow rate?’. Fortunately, with the constructed KG incorporating structural nodes and our designed traversal algorithm retrieving structural contexts, our proposed method achieves 67% Struct-EM.
@@ 7 Conclusion
Ethics Statement
Due to page limitation, the supplementary material and reproducing details are publically available at https://github.com/YuWVandy/KG-LLM-MDQA/blob/main/AAAI24˙KGPrompting˙Supplementary.pdf.
@@ Acknowledgements
This research is supported by Adobe Research and the National Science Foundation (NSF) under grant number IIS2239881.